### v1

@@ 1 Introduction
The emerging field of explainable artificial intelligence (XAI) has been addressing issues related to understanding the behavior of complex models, including LLMs. Many XAI methods focus on the simpler problem of local explanation, which attempts to explain specific model predictions rather than general model behavior. One popular local XAI methodology is feature attribution, which explains a model output by scoring the relative importance of the input features. For RAG systems, context attribution can assign scores to each piece of information retrieved by an LLM system at inference time [11]. However, feature attribution is not actionable (e.g., to achieve recourse).
Counterfactual explanations address this problem, explaining a model’s output by citing a minimal input perturbation that would induce an alternative output [37]. In RAG, a counterfactual explanation might reveal that the LLM would have generated a different answer if some piece of information were not provided or retrieved at runtime [30]. However, there may be many minimal counterfactuals for a given prediction, which may overwhelm end users. Explanations using if-then rules mitigate these problems by summarizing patterns observed across a model’s predictions [33, 26, 15], but have not yet been considered in the context of LLMs with RAG.
We fill this research gap and present a rule framework for explaining the output of black-box RAG systems using rules that formulate input predicates based on the presence of RAG sources provided to the LLM. Let us illustrate using a healthcare example. Suppose that a clinic is using a RAG system connected to a medical database, and a medical research assistant working at the clinic asks the LLM, “What treatment is most effective for Long COVID fatigue?” Here, a rule could reveal that “Whenever ‘Document 123’ is included in the RAG sources, the LLM always recommends an unsafe Long COVID treatment.” (A rule’s antecedent could likewise be when a given document is omitted from the RAG sources.) This rule helps curators of the medical database flag documents such as Document 123 that, unbeknownst to them, advocate for unsafe treatments. In this case, Document 123 may advocate for a treatment that was thought to be effective until recently, while newer studies have shown it is both ineffective and unsafe for certain demographics. Our framework empowers developers, auditors, policy makers, and other end users to seek rules for any identifiable output condition, such as responses that are incorrect, contain misinformation, or reflect negative sentiment [19].
@@ 2 Related Work
In the XAI literature, many methods have been proposed to explain individual model predictions, collectively known as local explanations. Many local methods adopt an approach known as feature attribution, exemplified by methods such as LIME [32] and SHAP [24], which explain a model by scoring input features according to their relative influence on the model’s output. Though intuitive, these explanations are not directly actionable, and have been shown to be potentially unreliable, misleading, and manipulable [22, 34, 21]. In contrast, counterfactual explanations explain a model by perturbing the input features in a minimal way to change the original output [37, 36]; e.g., a declined loan application would have been approved if the applicant’s income were 20 percent higher. Counterfactual perturbations aim to produce actionable insights that facilitate recourse, but many valid counterfactuals may exist, potentially contradicting each other and overwhelming users. Some methods therefore seek to sample diverse counterfactual perturbations or prioritize those deemed more feasible [27, 35].
@@
At the cost of increased computational complexity, rules mitigate the above problems by summarizing patterns observed across a model’s predictions [33, 26, 15, 31]; e.g., “if employed = no, then loan application = rejected,” regardless of the values of the other features. Recent works have established that rules essentially subsume counterfactuals, forming a duality, since both make assertions over a model’s input features [17, 16]. Rule-based explanations have been defined as conjunctions of predicates over the features of an input, for which any matching input always leads to the model producing the same output [31]. Several methods, such as Anchors [33], have applied similar definitions for the purposes of XAI. In contrast to our work, their rules are approximate (i.e., hold with high confidence), not exact, and do not provide an interface to instantiate their rules’ input predicates for modern LLMs or RAG systems.
@@
These rule formulations build implicitly upon association rule formulations established in the rule mining literature [2, 3, 6]. These formulations gave rise to many influential rule mining algorithms that search for frequent itemsets and valid association rules, most notably the Apriori algorithm [3]. The rule mining methods we present in this work take algorithmic inspiration from Apriori, but have several key modifications for a fundamentally different problem: explaining AI model predictions. While Apriori iterates over observed itemsets, we instead iterate over the powerset of possible model inputs (features), eliminating any concept of support (see Definition 2) in rule validity. This enables us to seek sets that satisfy arbitrary predicates over model outputs, as opposed to those observed with some minimum frequency, as in the original Apriori algorithm. Our algorithms also leverage the concept of monotonicity (i.e., the Apriori property) to prune the search space, but we propose a new definition that operates on model output predicates rather than itemset support.
@@
Naturally, many recent XAI works have focused on explaining LLMs, RAG, and other related phenomena, adapting general XAI methodologies as needed. Global explanation disciplines such as mechanistic interpretability attempt to identify low-level circuits that form among LLM neurons [9, 14, 38]. However, these insights are intended primarily for scientific (i.e., expert) understanding. One intuitive explanation format for LLMs and RAG systems are citations, which can be generated at inference time using specialized training and fine-tuning methods [23, 5] or post-hoc using various citation recommendation strategies [13, 28, 25]. However, these methods are subject to the critical limitation of hallucination [4], as their predictive approach to generating citations can be highly inaccurate, even with state-of-the-art models [29, 10, 7, 12, 18]. Meanwhile, recent works have adapted feature attribution [11] and counterfactual methods [30] to explain source importance in RAG, assigning importance weights to the retrieved sources or generating counterfactuals by removing some sources to change the LLM’s output, respectively. In contrast to our rule-based approach, these recent methods are prone to the same limitations as their general methodologies, and may fail to produce concise and actionable insights that are robust and guaranteed.
@@ 3.1 Rule Design
In the data mining literature, a rule expresses a conditional “if-then” relationship where satisfying the “if” part (the antecedent) implies that the “then” part (the consequent) is also satisfied [6]. One intuitive adaptation for rules describing a prediction $y=M(x)$ might define the antecedent in terms of the input $x$ to the model $M$, and the consequent in terms of the resulting output $y$. This would yield rules of the form “if $x$ then yy”. We adopt this general orientation due to its clear alignment with the primary goal of local XAI: understanding the effect of model inputs on model outputs. This orientation has also been adopted in recent XAI works, wherein rules explain model predictions by uncovering significant patterns in said predictions. In this XAI setting, rules are better known as rule-based explanations.
@@
A rule-based explanation is a conjunction of predicates over the features of some model $M$, where for any input $x^{*}$ satisfying all predicates, the model always generates some target output $y^{*}$ [17, 31].
@@
In settings that observe such instances, rules are considered more informative when they have higher support (i.e., the rule is more likely to implicate an instance) and higher confidence (i.e., the rule is more likely to hold). In this work, we do not seek rules that characterize observed data, and thus do not presume a setting wherein a set of observed instances is available. Instead, we seek rules that characterize the full model input space, and therefore operate over the set of all possible antecedents. It follows that the classical notion of support does not apply in our setting, though it is nonetheless equal to the reciprocal of the number of possible antecedents. Furthermore, we require that all valid rules hold with certainty (i.e., confidence of 100%), but note that future work could build upon existing rule methods that guarantee minimum confidence levels [33].
@@
Drawing inspiration from XAI literature, we contend that feature ablation (i.e., removing text segments) is a suitable perturbation strategy for explaining language models, and thus formalize antecedent predicates based on feature presence. In the XAI literature, ablation is a well-founded strategy for inferring feature importance [20], and has seen applications in works that explain language models [30, 11] and models generally [24]. We assume that the user partitions the input $x$ into two components $x=(\mathbf{s},\mathbf{c})$, where $\mathbf{s}$ represents some feature subset of interest $\mathbf{\mathbf{s}}=\{s_{1},s_{2},\dots,s_{n}\}$ derived from the input (e.g., RAG document sources). In our rule formulation, the antecedent will be defined in terms of the presence (or similarly, the absence) of inputs in $\mathbf{s}$, while the remaining inputs in $\mathbf{c}$ are ignored. By separating $\mathbf{s}$ and $\mathbf{c}$ (i.e., $y=M(\mathbf{s^{\prime}},\mathbf{c})$), we enable the user to focus their rules and explanations on small, interesting subsets of the input feature space, while reducing the complexity of our rule mining algorithms.
@@
Consequent: We contend that the ideal choice of consequent is ultimately dictated by the user and their motivation for seeking rules. Therefore, the consequent of our rules asserts only that the output meets some target condition defined explicitly by the user. In essence, the target condition is an arbitrary output predicate that could test for any identifiable property (e.g., presence of misinformation, validity of in-text citations). We expect the user to define an output predicate function $O:\mathcal{Y}\rightarrow\{0,1\}$ that returns 1 when a model output $y\in\mathcal{Y}$ meets these conditions of interest. By parameterizing the consequent, we empower the user to seek rules that implicate not just the presence of a specific output (as is common in many XAI and rule-based works), but more sophisticated output characteristics (e.g., presence of hallucination, sentiment).
@@
To define an output predicate function $O$, the user must effectively implement a classifier that inspects an arbitrary LLM output and judges whether it meets the conditions of interest. A simple predicate might compare the answer for equality against some target answer, returning 1 if the LLM answer is equal to the target answer (and 0 otherwise). A more specific predicate might analyze the LLM answer for correctness; however, this would only be possible in settings where the correct answer is known upfront (e.g., benchmarking). For nuanced semantic assertions (e.g., judging citation validity), or those that must be robust to verbose or unstructured LLM outputs (e.g., spotting inappropriate content), a predicate function could adopt an “LLM-as-a-judge” implementation. Given detailed instructions and necessary context (e.g., ground truth information), an LLM can be a highly effective proxy for human judgment of LLM responses [8, 40]; however, stochasticity introduced during the decoding process (e.g., token sampling) can compromise reproducibility.
@@ 3.2 Rule Formulation
A retention rule identifies a subset of the input set that, when retained, always satisfies the output predicate. Similarly, an omission rules identifies a subset that, when omitted, always satisfies the output predicate. Inclusion and omission rules will often correlate with outputs that satisfy reciprocal (i.e., complementary) output predicates. Therefore, users will often use two separate, reciprocal output predicates when explaining a single input (e.g., $O_{correct}$ and $O_{incorrect}$). Generally, both rule types identify critical inputs in $\mathbf{s}$ whose presence is necessary to ensure some condition holds.
@@ 3.3 Rule Validity
By definition, all ancestors must be validated before validating a descendant. A rule mining algorithm must first validate those rules that retain or omit the most inputs (i.e., fully recurse) before unwinding the call stack to consider rules that retain or omit fewer inputs. Since our rules are expressed in terms of a set of features $\mathbf{s^{\prime}}$, this relationship can be characterized as set subsumption. Ancestor rules are subsumed by descendant rules, and are therefore redundant to the end user.
@@
A rule $r_{1}$ subsumes another rule $r_{2}$ iff every predicate in $r_{2}$ also appears in $r_{1}$.
@@
In this work, we seek minimal rules, but operate under the understanding that less-minimal rules must be validated first. We acknowledge an inherent trade-off between rule minimality and algorithmic efficiency, as finding more minimal (higher coverage) rules ultimately requires validating more rules. Our algorithms (introduced in Section 4) ultimately prioritize minimality, exhausting all rule candidates and guaranteeing 100% confidence. However, to eliminate redundancy, our explanations discard any non-minimal rules before being returned to the user.
@@
The full hierarchy of a given rule is a subset of the power set $\mathcal{P}(\mathbf{s})$. The power set, which contains all candidate rules that implicate $\mathbf{s}$, forms a data structure known as a lattice [1], illustrated in Figure 2. Each of the $2^{|\mathbf{s}|}$ subsets in $\mathcal{P}(\mathbf{s})$ is a rule candidate; each node in the lattice is a candidate, and each edge delineates a subset / superset relationship between two candidates. The $\mathcal{P}(\mathbf{s})$ lattice constitutes the full search space for any rule mining algorithm under our formulation, and it contains $2^{|\mathbf{s}|}$ candidate rules. In the following subsections, we present two rule mining algorithms that efficiently navigate this lattice in search of all valid rules.
@@ 4 Rule Mining Algorithms
In this section, we introduce two search algorithms capable of mining our rule-based explanations, each with different advantages. First, we introduce the Mono Rule Miner, which can mine either retention or omission rules in a single pass. Second, we introduce the Dual Rule Miner, which can mine both rule types simultaneously in a single pass. Mono is more efficient in settings where a single rule type is sufficient, and can exploit greater search space pruning. For settings where patterns of both types are sought and compared, Dual is often more efficient than executing two Mono passes.

### v2

@@ 1 Introduction
The emerging field of explainable artificial intelligence (XAI) has been addressing issues related to understanding the behavior of complex models, including LLMs. Many XAI methods focus on the simpler problem of local explanation, which attempts to explain specific model predictions rather than general model behavior. One popular local XAI methodology is feature attribution, which explains a model output by scoring the relative importance of the input features. For RAG systems, context attribution can assign scores to each piece of information retrieved by an LLM system at inference time [contextcite]. However, feature attribution is not actionable (e.g., to achieve recourse).
Counterfactual explanations address this problem, explaining a model’s output by citing a minimal input perturbation that would induce an alternative output [wachter2017counterfactual]. In RAG, a counterfactual explanation might reveal that the LLM would have generated a different answer if some piece of information were not provided or retrieved at runtime [RAGE]. However, there may be many minimal counterfactuals for a given prediction, which may overwhelm end users. Explanations using if-then rules mitigate these problems by summarizing patterns observed across a model’s predictions [anchors, MACHA2022101209, lore], but have not yet been considered in the context of LLMs with RAG.
We fill this research gap and present a rule framework for explaining the output of black-box RAG systems using rules that formulate input predicates based on the presence of RAG sources provided to the LLM. Let us illustrate using a healthcare example. Suppose that a clinic is using a RAG system connected to a medical database, and a medical research assistant working at the clinic asks the LLM, “What treatment is most effective for Long COVID fatigue?” Here, a rule could reveal that “Whenever ‘Document 123’ is included in the RAG sources, the LLM always recommends an unsafe Long COVID treatment.” (A rule’s antecedent could likewise be when a given document is omitted from the RAG sources.) This rule helps curators of the medical database flag documents such as Document 123 that, unbeknownst to them, advocate for unsafe treatments. In this case, Document 123 may advocate for a treatment that was thought to be effective until recently, while newer studies have shown it is both ineffective and unsafe for certain demographics. Our framework empowers developers, auditors, policy makers, and other end users to seek rules for any identifiable output condition, such as responses that are incorrect, contain misinformation, or reflect negative sentiment [han2024medical].
@@ 2 Related Work
In the XAI literature, many methods have been proposed to explain individual model predictions, collectively known as local explanations. Many local methods adopt an approach known as feature attribution, exemplified by methods such as LIME [ribeiro2016should] and SHAP [shap], which explain a model by scoring input features according to their relative influence on the model’s output. Though intuitive, these explanations are not directly actionable, and have been shown to be potentially unreliable, misleading, and manipulable [Kindermans2019, advlime:aies20, shap-edu-problems]. In contrast, counterfactual explanations explain a model by perturbing the input features in a minimal way to change the original output [wachter2017counterfactual, verma2024counterfactual]; e.g., a declined loan application would have been approved if the applicant’s income were 20 percent higher. Counterfactual perturbations aim to produce actionable insights that facilitate recourse, but many valid counterfactuals may exist, potentially contradicting each other and overwhelming users. Some methods therefore seek to sample diverse counterfactual perturbations or prioritize those deemed more feasible [mothilal2020dice, NEURIPS2020_c2ba1bc5].
@@
At the cost of increased computational complexity, rules mitigate the above problems by summarizing patterns observed across a model’s predictions [anchors, MACHA2022101209, lore, rudinrules]; e.g., “if employed = no, then loan application = rejected,” regardless of the values of the other features. Recent works have established that rules essentially subsume counterfactuals, forming a duality, since both make assertions over a model’s input features [geng2022computing, guidotti2024stable]. Rule-based explanations have been defined as conjunctions of predicates over the features of an input, for which any matching input always leads to the model producing the same output [rudinrules]. Several methods, such as Anchors [anchors], have applied similar definitions for the purposes of XAI. In contrast to our work, their rules are approximate (i.e., hold with high confidence), not exact, and do not provide an interface to instantiate their rules’ input predicates for modern LLMs or RAG systems.
@@
These rule formulations build implicitly upon association rule formulations established in the rule mining literature [rules-1993, apriori, most-interesting-rules]. These formulations gave rise to many influential rule mining algorithms that search for frequent itemsets and valid association rules, most notably the Apriori algorithm [apriori] and more-efficient successors like Eclat [eclat] and FP-growth [fpgrowth]. The rule mining methods we present in this work take algorithmic inspiration from Apriori, but have several key modifications for a fundamentally different problem: explaining AI model predictions. While Apriori iterates over observed itemsets, we instead iterate over the powerset of possible model inputs (features), eliminating any concept of support (see Definition 2) in rule validity. This enables us to seek sets that satisfy arbitrary predicates over model outputs, as opposed to those observed with some minimum frequency, as in the original Apriori algorithm. Our algorithms also leverage the concept of monotonicity (i.e., the Apriori property) to prune the search space, but we propose a new definition that operates on model output predicates rather than itemset support.
@@
Naturally, many recent XAI works have focused on explaining LLMs, RAG, and other related phenomena, adapting general XAI methodologies as needed. Global explanation disciplines such as mechanistic interpretability attempt to identify low-level circuits that form among LLM neurons [autocircuits, NEURIPS2024_a8f7d43a, NEURIPS2024_ad217e0c]. However, these insights are intended primarily for scientific (i.e., expert) understanding. One intuitive explanation format for LLMs and RAG systems are citations, which can be generated at inference time using specialized training and fine-tuning methods [sourceaware2024, aly-etal-2024-learning] or post-hoc using various citation recommendation strategies [farber2020citation, maheshwari-etal-2025-citefix, ceg]. However, these methods are subject to the critical limitation of hallucination [agrawal-etal-2024-language], as their predictive approach to generating citations can be highly inaccurate, even with state-of-the-art models [citeme, cao-wang-2024-verifiable, byun-etal-2024-reference, day2023preliminary, alce]. Meanwhile, recent works have adapted feature attribution [contextcite] and counterfactual methods [RAGE] to explain source importance in RAG, assigning importance weights to the retrieved sources or generating counterfactuals by removing some sources to change the LLM’s output, respectively. In contrast to our rule-based approach, these recent methods are prone to the same limitations as their general methodologies, and may fail to produce concise and actionable insights that are robust and guaranteed.
@@ 3.1 Rule Design
In the data mining literature, a rule expresses a conditional “if-then” relationship where satisfying the “if” part (the antecedent) implies that the “then” part (the consequent) is also satisfied [most-interesting-rules]. One intuitive adaptation for rules describing a prediction $y=M(x)$ might define the antecedent in terms of the input $x$ to the model $M$, and the consequent in terms of the resulting output $y$. This would yield rules of the form “if $x$ then yy”. We adopt this general orientation due to its clear alignment with the primary goal of local XAI: understanding the effect of model inputs on model outputs. This orientation has also been adopted in recent XAI works, wherein rules explain model predictions by uncovering significant patterns in said predictions. In this XAI setting, rules are better known as rule-based explanations.
@@
A rule-based explanation is a conjunction of predicates over the features of some model $M$, where for any input $x^{*}$ satisfying all predicates, the model always generates some target output $y^{*}$ [geng2022computing, rudinrules].
@@
In settings that observe such instances, rules are considered more informative when they have higher support (i.e., the rule is more likely to implicate an instance) and higher confidence (i.e., the rule is more likely to hold). In this work, we do not seek rules that characterize observed data, and thus do not presume a setting wherein a set of observed instances is available. Instead, we seek rules that characterize the full model input space, and therefore operate over the set of all possible antecedents. It follows that the classical notion of support does not apply in our setting, though it is nonetheless equal to the reciprocal of the number of possible antecedents. Furthermore, we require that all valid rules hold with certainty (i.e., confidence of 100%), but note that future work could build upon existing rule methods that guarantee minimum confidence levels [anchors].
@@
Drawing inspiration from XAI literature, we contend that feature ablation (i.e., removing text segments) is a suitable perturbation strategy for explaining language models, and thus formalize antecedent predicates based on feature presence. In the XAI literature, ablation is a well-founded strategy for inferring feature importance [hameed:basedxai:2022], and has seen applications in works that explain language models [RAGE, contextcite] and models generally [shap]. We assume that the user partitions the input $x$ into two components $x=(\mathbf{s},\mathbf{c})$, where $\mathbf{s}$ represents some feature subset of interest $\mathbf{\mathbf{s}}=\{s_{1},s_{2},\dots,s_{n}\}$ derived from the input (e.g., RAG document sources). In our rule formulation, the antecedent will be defined in terms of the presence (or similarly, the absence) of inputs in $\mathbf{s}$, while the remaining inputs in $\mathbf{c}$ are ignored. By separating $\mathbf{s}$ and $\mathbf{c}$ (i.e., $y=M(\mathbf{s^{\prime}},\mathbf{c})$), we enable the user to focus their rules and explanations on small, interesting subsets of the input feature space, while reducing the complexity of our rule mining algorithms.
@@
Consequent: We contend that the ideal choice of consequent is ultimately dictated by the user and their motivation for seeking rules. Therefore, the consequent of our rules asserts only that the output meets some target condition defined explicitly by the user. In essence, the target condition is an arbitrary output predicate that could test for any identifiable property (e.g., presence of misinformation, validity of in-text citations). We expect the user to define an output predicate function $O:\mathcal{Y}\rightarrow\{0,1\}$ that returns 1 when a model output $y\in\mathcal{Y}$ meets these conditions of interest. By parameterizing the consequent, we empower the user to seek rules that implicate not just the presence of a specific output (as is common in many XAI and rule-based works), but more sophisticated output characteristics (e.g., presence of hallucination, bias, sentiment).
@@
To define an output predicate function $O$, the user must effectively implement a classifier that inspects an arbitrary LLM output and judges whether it meets the conditions of interest. A simple predicate might compare the answer for equality against some target answer, returning 1 if the LLM answer is equal to the target answer (and 0 otherwise). A more specific predicate might analyze the LLM answer for correctness; however, this would only be possible in settings where the correct answer is known upfront (e.g., benchmarking). For nuanced semantic assertions (e.g., judging citation validity), or those that must be robust to verbose or unstructured LLM outputs (e.g., spotting inappropriate content), a predicate function could adopt an “LLM-as-a-judge” implementation. Given detailed instructions and necessary context (e.g., ground truth information), an LLM can be a highly effective proxy for human judgment of LLM responses [chiang-lee-2023-large, judging-llm-judge]; however, stochasticity introduced during the decoding process (e.g., token sampling) can compromise reproducibility.
@@ 3.2 Rule Formulation
A retention rule identifies a subset of the input set that, when retained, always satisfies the output predicate. Similarly, an omission rules identifies a subset that, when omitted, always satisfies the output predicate. Retention and omission rules will often correlate with outputs that satisfy reciprocal (i.e., complementary) output predicates. Therefore, users will often use two separate, reciprocal output predicates when explaining a single input (e.g., $O_{correct}$ and $O_{incorrect}$). Generally, both rule types identify critical inputs in $\mathbf{s}$ whose presence is necessary to ensure some condition holds.
@@ 3.3 Rule Validity
By definition, all ancestors must be validated before validating a descendant. A rule mining algorithm must first validate those rules that retain or omit the most inputs (i.e., fully recurse) before unwinding the call stack to consider rules that retain or omit fewer inputs. Since a rule implicates all ancestor rules whose representative subset is a superset of its own, rules representing smaller subsets will implicate more rules than those representing larger, meaning that descendant rules have greater coverage than ancestor rules. Under this interpretation, descendant rules can be seen as subsuming ancestor rules as the latter have less coverage (implicate fewer feature combinations).
@@
A rule $r_{1}$ subsumes another rule $r_{2}$ iff every predicate in $r_{1}$ also appears in $r_{2}$.
@@
In this work, we seek minimal rules, but operate under the understanding that less-minimal rules must be validated first. We acknowledge an inherent trade-off between rule minimality and algorithmic efficiency, as finding more minimal (higher coverage) rules ultimately requires validating more rules. Our algorithms (introduced in Section 4) ultimately prioritize minimality, exhausting all rule candidates and guaranteeing 100% confidence. However, to eliminate redundancy, non-minimal rules can be discarded before being returned to the user.
@@
The full hierarchy of a given rule is a subset of the power set $\mathcal{P}(\mathbf{s})$. The power set, which contains all candidate rules that implicate $\mathbf{s}$, forms a data structure known as a lattice [abiteboul1995foundations], illustrated in Figure 2. Each of the $2^{|\mathbf{s}|}$ subsets in $\mathcal{P}(\mathbf{s})$ is a rule candidate; each node in the lattice is a candidate, and each edge delineates a subset / superset relationship between two candidates. The $\mathcal{P}(\mathbf{s})$ lattice constitutes the full search space for any rule mining algorithm under our formulation, and it contains $2^{|\mathbf{s}|}$ candidate rules. In the following subsections, we present two rule mining algorithms that efficiently navigate this lattice in search of all valid rules.
@@ 4 Rule Mining Algorithms
In this section, we introduce two search algorithms [^1] Algorithm implementations are available at https://github.com/joelrorseth/rag-llm-rule-miner. capable of mining our rule-based explanations, each with different advantages. First, we introduce the Mono Rule Miner, which can mine either retention or omission rules in a single pass. Second, we introduce the Dual Rule Miner, which can mine both rule types simultaneously in a single pass. Mono is more efficient in settings where a single rule type is sufficient, and can exploit greater search space pruning. For settings where patterns of both types are sought and compared, Dual is often more efficient than executing two Mono passes.